OpenStudy (anonymous):

I've tried this for an Hour, and can't get it. A metallurgist has an alloy with 15% TITANIUM and an alloy with 30% TITANIUM. He needs 100 GRAMS of an alloy with 21% TITANIUM. How much of each alloy should be mixed to attain the 100 Grams of Alloy with 21% Titanium?

OpenStudy (anonymous):

I know i need 2 equations and I have to use elimation to find the answers.

OpenStudy (mertsj):

x= amount of 15% alloy 100-x= amount of 30% alloy to use .15x= amount of titanium in x grams of 15% alloy .30(100-x)= amount of titanium in 100-x grams of 30 % alloy .21(100) = amount of titanium in 100 grams of 21% alloy .15x+.30(100-x)=.21(100) Or in other words, initial titanium + added titanium = final titanium

OpenStudy (mertsj):

\[.15x+.3(100-x)=.21(100)\] \[.15x+30-.3x=21\] \[-.15x=-9\] \[x=60\] Use 60 g of15% alloy and 40 g of 30% alloy And NO. You don't need two equations. Although you could do it with 2 and they would be: x+y=100 .15x+.3y=.21(100)

OpenStudy (anonymous):

Thanks, I was setting it up wrong at the begining. I understand now. Thank you

OpenStudy (mertsj):

yw